OpenStudy (anonymous):

Could you help me solve these? |2x - 3| = -17 |5x + 3| = 12

OpenStudy (turingtest):

for the first one, what happens when you take the absolute value of a negative number?

OpenStudy (anonymous):

Doesnt it turn into a positive

OpenStudy (turingtest):

correct, so how can taking the absolute value of the left side leave us with -17 ?

OpenStudy (anonymous):

I dont know?

OpenStudy (turingtest):

the answer is you can't; it's impossible hence the answer to the first problem is...?

OpenStudy (anonymous):

NO solution?

OpenStudy (turingtest):

correct

OpenStudy (turingtest):

for the next one, the fact that taking the absolute value of the left gives us 12 means that the left side is one of which two possible numbers?

OpenStudy (anonymous):

It would be a negative number

OpenStudy (anonymous):

Wouldnt you divide 5 from both sides then multiply the number by 3 and negative 3?

OpenStudy (turingtest):

no, what two numbers have an absolute value of 12?

OpenStudy (anonymous):

-12 and 12

OpenStudy (turingtest):

great, so solve the left side twice, once for each possibility

OpenStudy (anonymous):

I dont know how to do that

OpenStudy (anonymous):

You want to get x alone dont you?

OpenStudy (turingtest):

|5x + 3| = 12 means that either 5x + 3 = 12 or 5x + 3 = -12

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

So you would divide 5 from both sides

OpenStudy (turingtest):

i would do that last

OpenStudy (anonymous):

So what should i do first?

OpenStudy (turingtest):

deal with the addition

OpenStudy (anonymous):

Subtract 3?

OpenStudy (turingtest):

yep

OpenStudy (anonymous):

Okay so for the equation with positive 12 it is 1.8

OpenStudy (turingtest):

sounds right

OpenStudy (anonymous):

Then for the second equation it is -3

OpenStudy (turingtest):

yep :)

OpenStudy (anonymous):

and thats all you need to do?

OpenStudy (turingtest):

that's it

OpenStudy (turingtest):

you would write the answer as\[x=\{-3,1.8\}\]

OpenStudy (anonymous):

So equation one had no solutions and equation two and 2 solutions?

OpenStudy (turingtest):

that's correct

OpenStudy (anonymous):

@TuringTest I have another questions about absolute values.. concerning the same thing we just went over

OpenStudy (anonymous):

-3|x - 3| = -6 with this equation right here.. What would i do with the negative 3 thats outside of the lines?

OpenStudy (anonymous):

Would i add that up in the end or what?

OpenStudy (turingtest):

no, you need to deal with that first treat |x-3| as the thing you need to find. let's say y=|x-3|. the problem is then -3y=-6

OpenStudy (turingtest):

simplify this to get y=2 and now substitute back in |x-3|=2